Bank 1 qoutes an annual rate of 11.2 % with monthlly compounding. Bank 2 uses semiannual compounding and quotes an annual rate of 11.4 %

Requirement 1:

Calculate the effective annual interest rate for each bank. (Enter your answers as a percent rounded to 2 decimal places (e.g., 32.16).)

EAR
Bank 1 %
Bank 2 %

Thank you!!
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

Bank 1:

EAR=(1+0.112/12)^12-1

=11.79%(Approx).

Bank 2:

EAR=(1+0.114/2)^2-1

=11.72%(Approx).